OpenStudy (anonymous):

@SandeepReddy Given the events below, determine which equation correctly calculates the probability of drawing two kings in a row from a standard 52-card deck, without replacement. Event A: The first card drawn is a king. Event B: The second card drawn is a king. A. P(A n B) = P(A) * P(B|A) B. P(A n B) = P(A) * P(B) C. P(A n B) = P(A) * P(A|B) D. P(A n B) = P(B) * P(B|A)

OpenStudy (anonymous):

Sorry for the delay Have you ever played or seen Cards?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

so how many kings wil be in a deck?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

now probability from basic definition \[P = \frac{ Favourable._.events }{ Total_.Number_.of_.Events }\]

OpenStudy (anonymous):

for example, lets calculate the probability of getting even number when a 6 sided dice is thrown Now in this case : favourable events are 2,4,6 total 3 and total number of possible events are 1,2,3,4,5,6 total 6 so probability of getting even number when a 6 sided dice is 3/6 = 0.5

OpenStudy (anonymous):

@ashleigh0519 interrupt whenever u feel that something is not clear to u

OpenStudy (anonymous):

to continue

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Having said about calculating the probability then, thats all u need to know to solve this problem :)

OpenStudy (anonymous):

Now lets calculate the probability of event A: favourable cases = 4 total events = 52 probability P(A) = 4/52 = 1/13

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

Now coming to the probability of B:

OpenStudy (anonymous):

Now that already one King has gone in event A and furthermore, they said "without replacement" so Can u tell the favourable and total events for B?

OpenStudy (anonymous):

P(B) = 3/52

OpenStudy (anonymous):

You have at first total 52 cards, so total events in first event A is 52. Now you took out a King from a deck of 52 cards in even A so how many cards are there in total for event B?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

how 50?

OpenStudy (anonymous):

wait 51

OpenStudy (anonymous):